OpenStudy (anonymous):

integral of sqrt(36+x^2)dx Can anyone help me with this? I used trig identities and somehow got to integral of 36 sec^3(θ) dθ but I'm not sure if this is right or what my next step is..

OpenStudy (anonymous):

What is an integral?

OpenStudy (anonymous):

Use a substitution \(x=6\tan(z)\), the integral will be easy with this substitution. After you evaluate the integral in terms of \(z\), substitute back for \(x\).

OpenStudy (anonymous):

I did that and got \[36\sqrt{1+\tan^2\theta} (\sec^2\theta) d \theta\] but I'm not sure where to go from here. Another substitution?

OpenStudy (anonymous):

That expression can be simplified into \(36\sec^3\theta\). To evaluate this integral, use integration by parts with \(u=\sec\theta \) and \(dv=\sec^2 xdx\).

OpenStudy (anonymous):

Try to do it and let me know if you're stuck at any point.

OpenStudy (anonymous):

okay so I have this \[36(\sec \theta \tan \theta - \int\limits_{}^{} \tan^2 \theta \sec \theta d \theta\] but I'm stuck at the second part of the integral I made tan^(θ) = 1-sec^2(θ) and set u = secθ, dv = 1-sec^2(θ) Is this correct?

OpenStudy (anonymous):

Let our integral be \(I\). Then by integration by parts \[I=\sec\theta \tan\theta -\int \tan^2\theta \sec\theta d\theta=\sec\theta\tan\theta-\int\sec^3\theta d\theta+\int \sec\theta d\theta. \] So, \(I=\sec\theta\tan\theta-I+\int \sec\theta d\theta \implies \large { 2I=\sec\theta \tan\theta+\int\sec\theta d\theta}. \) Evaluate \(\int\sec\theta\) and solve for \(\bf I\).

OpenStudy (anonymous):

Don't forget to multiply by \(36\) at the end, since the original integral is \(36 I\).

OpenStudy (anonymous):

I got it now, thanks :)